# SynLlama: Generating Synthesizable Molecules and Their Analogs with Large Language Models

### v4

@@ 1 Introduction
Chemical space is enormous, built up via the exponential rise in functional group combinatorics that define an increasing diverse set of molecules. Traditional approaches that design synthetic pathways of unseen molecules under well-controlled laboratory conditions have been made possible by decades of research in synthetic chemistry, as well as mechanistic studies of key reaction steps and reaction classification1, 2. Using the wealth of data accumulated in libraries of chemical reactions, expert systems3, 4 have been developed that deploy this knowledge to construct multi-step pathways to specified end-products. Such methods have become a key tool for the bench chemist, as illustrated by the Chematica software and and its follow on commercial product now known as Synthia4.
@@
With recent advances in artificial intelligence and deep learning, generative models have begun to contribute to enumerating molecules at the stoichiometric scale. After training on databases containing various small molecules representations5, 6, 7, 8, 9, string-based 1D generative models and structure-aware 3D de novo methods have paved the way for quick exploration of greater swaths of unseen chemical space10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27. However, even with their exceptional generative capabilities, these models still face one major challenge: their proposed de novo molecules lack practical guarantees of synthesizability, which limits their utility in practice28, 29, 30. For generative approaches in drug and materials discovery to fulfill their potential, ensuring synthetic feasibility is essential to bridge the gap between in silico molecule design and the realistic applicability of computationally generated molecules.
@@
In recognition of this dissonance, efforts have been made to address the problem of poor synthesizability of de-novo-generated molecules. One line of research focuses on integrating empirical or deep-learning scoring functions31, 32, 33, 34, 35, 36, 37, such as the synthetic accessibility (SA) score 31 and DeepSA score35, into the objective functions of learning algorithms. However, optimizing only the synthesizability score can still lead to the generation of unsynthesizable molecules because the scoring functions rely on identifying common fragments or reactive centers in molecules38. In addition, they often assign bad scores to complex yet synthesizable molecules that require multi-step synthetic pathways, causing generative models to miss viable candidates 39. Others have proposed improving synthesizability by building molecules from common molecular fragments40, 41, 42, 43, but they still don’t guarantee synthesis as these methods do not explicitly consider the reaction pathways to build the molecular candidates. Another line of research integrates the explicit use of computer-assisted synthetic planning (CASP) software44, 45 into the optimization46. However, the computational overhead is significant and the quality of the optimized molecules can be variable.47.
@@
Alternatively, proposing synthesizable molecular candidates using commercially available building blocks and commonly known organic reaction templates48, 44 offers better synthetic tractability over simple molecule scoring. Importantly, this strategy is appealing to bench and medicinal chemists, since it offers actionable synthesis pathways for them to examine, refine, and execute. Some recent models in this direction apply rule-based synthesis and optimization on building blocks or entire synthetic pathways to generate novel molecules with desired chemical properties4, 49, 50, 51, 52, 53. Other models condition on input molecules to propose synthetic pathways using commercially available building blocks and well-validated reaction templates for either full construction of the target molecule or the generation of structurally similar analogs in a forward synthesis manner within the predefined chemical search space.54, 55, 56 For example, SynNet54 constructs synthetic trees via Markov Decision Processes (MDPs) and uses multilayer perceptrons to choose the next action space. More recent models such as ChemProjector55 and Synformer56 use transformers to decode for the next action space and have achieved good empirical performances for target and analog molecule reconstruction.
@@
A compelling alternative is the use of Large Language Models (LLMs) due to their foundational nature and adaptability to downstream tasks. 57 LLMs inherently possess extensive chemical knowledge, and recent advancements have focused on extracting and applying this knowledge for predictive and optimization tasks using natural language guidance58, 59, 60, 61. Furthermore, after fine-tuning, LLM models can perform as good or better than chemical language models trained solely on chemical representations, all while requiring less data.27 The efficiency and unexpected performance gains from fine-tuning LLMs thus motivates us to explore their potential in more complex tasks, such as synthesis planning, which could pave the way for new chemical discoveries.
@@
Herein, we present SynLlama, an LLM-based tool built on the open-source Llama-3.1-8B and Llama-3.2-1B foundation models  62 to deduce synthetic routes for target molecules or structurally related analogs. Specifically, the LLM component of SynLlama operates as a constrained retrosynthesis module that breaks input molecules into building blocks (BBs) via well-validated (RXN) sequences, and the reconstruction module searches commercially available BBs based on LLM predictions and builds up molecules within a diverse yet synthesizable chemical space. As an illustration of utility, SynLlama demonstrates competitive performance in key tasks for drug discovery, including synthesis planning for target and analog molecules of pharmaceutical interest and expansion around existing molecular drug hits and leads. Moreover, because of its generative nature, the LLM component of SynLlama has the added ability to explore commercially available building blocks beyond the predefined synthetic space introduced during training - an ability that previous models lack. By integrating molecular design with synthetic feasibility, SynLlama represents a step forward in bridging computational chemistry with synthetic chemistry, providing chemists with actionable and experimentally accessible molecular candidates.
@@ 2 Methods
The SynLlama workflow, illustrated in Figure 1, is designed to generate synthesizable compounds within an expanded chemical space. When an input molecule passes through this workflow, it can either be fully reconstructed through valid synthetic pathways, or the workflow will produce a structurally similar yet synthesizable analog along with its synthesis route. To transform general-purpose LLMs, like the Llama 3 models 62, into expert models for synthetic pathways, we use three key components: 1) a reliable and diverse set of reaction data that covers a large synthesizable chemical space, 2) an efficient supervised fine-tuning (SFT) strategy to train a general-purpose LLM on these reaction data, and 3) a reconstruction algorithm that can convert the output of the fine-tuned LLM into valid synthesis routes, ensuring the proposed molecules lie within a commercially available chemical search space. These components are crucial for leveraging LLMs, which are known to perform well in diverse chemistry tasks 63, 64, to specialize in synthetic modeling.
@@ 2.1 Reaction Data for Training and Testing Sets
As illustrated in Figure 1(a), our defined chemical space for training consists of molecules that can be synthesized in at most five steps with Enamine building blocks9 (BBs) and 2 sets of well-validated common organic reactions (RXNs). To define the training and testing BB data, we apply a time split whereby all Enamine BBs from the August 2024 release serve as the training BBs, and all new BBs from their February 2025 release that were not in the training set comprise the testing BBs. This procedure results in $\sim$230,000 BBs for training and $\sim$13,000 BBs for testing. Later in Section 3.1 we consider the target reconstruction for unseen molecules which are taken from the Enamine Diversity Set9 and ChEMBL dataset8.
We define two sets of reaction templates (RXN) which operate on the Enamine BBs. RXN 1 is formulated as a set of 91 reaction templates selected by Gao et al.54 from the works of Hartenfeller et al.65 and Button et al.66. RXN 2 is comprised of 115 reactions selected by Gao et al.56 that contains reactions used to create the Enamine REAL space9 plus some reactions from RXN 1. All reaction templates are accessible to both sets of BBs, thus defining the training and testing chemical spaces. As a result, there are $\sim$$10^{30}$ molecules within this space that can be represented by a synthesis path that comprises a sequence of BBs and RXNs.
@@ 2.2 Supervised Fine-tuning and Inference from SynLlama
To create the SynLlama model, we need to establish data generation protocols for supervised fine-tuning (SFT) of the Llama 3 models as schematically shown in Figure 1(c). When generating reaction data in text format, we choose to represent the BBs and intermediates along the synthetic pathway using SMILES67 strings, while RXNs are explicitly defined in the SMARTS68 format. These structured chemical notations are designed to enhance SynLlama’s ability to systematically identify and deconstruct bonds according to RXN templates, effectively dismantling input molecules into building-block-sized fragments.
@@ 2.3 SynLlama Model Benchmarks
In Table 1, all four trained SynLlama models are evaluated on both in-distribution training data and out-of-distribution testing and ChEMBL 8 data to assess the benchmarks outlined above. In the instruction-following benchmarks, most models exhibit strong adherence (over 90 percent) to the fine-tuned response structure across all datasets. This impressive performance indicates that fine-tuning effectively retains the specified output structure when trained with over 100,000 samples. Furthermore, all four models successfully memorized the provided RXN templates and selected the building blocks (BBs) from all predicted reactants over 99 percent of the time. This capability further enhances the coupling effectiveness of the downstream reconstruction algorithm with the SynLlama raw output, as it only requires information about reaction sequences and predicted building blocks.
In the reaction chemistry benchmarking results, a clearer trend emerges: models, regardless of their parameter size, show improved comprehension of reaction chemistry in all three datasets as the amount of training data increases. Notably, most models maintain their performance from training to testing data, but exhibit a greater decline in “Matched Reactants” and “Good Products” performance when generalizing to the ChEMBL data. The reason behind this is that the testing data are generated in the same manner as the training data but with a different set of building blocks, while the ChEMBL data occupies a different chemical space, as previously noted by Gao et al. 54. Despite the reductions in their performance for ChEMBL molecules, as shown in Supplementary Figure S2, SynLlama-8B-500k and SynLlama-1B-2M can still generate complete and valid syntheses over 50% of the time without any downstream processing. These results indicates that SynLlama’s raw results alone have potential utility for synthesis planning for unseen drug-like molecules.
@@ 2.4 Reconstruction from Predicted Retrosynthesis
Using the predicted sequence of RXNs and BBs from SynLlama responses, we can synthesize the proposed target molecule or close analogs by applying the predicted reaction templates to the BBs in the inferred order, as shown in black ink in Figure 1(d). In some cases, the predicted BBs match known Enamine BBs, ensuring that the resulting molecules remain within an established chemical space for synthesis. However, due to SynLlama’s generative nature, some predicted BBs are novel while still providing valid synthesis pathways, and we only report new BBs that can be purchased from other suppliers identified by Molport 69. Therefore, while SynLlama primarily produces molecules within the predefined chemical space using Enamine BBs, its output also offers an alternative strategy for molecule construction. We will revisit this point in the Results section.
When the input molecule cannot be fully reconstructed, we generate analogs by mapping the predicted BBs from SynLlama to known Enamine BBs, thereby sampling molecules from the well-defined Enamine chemical space. Under this scenario, we use nearest neighbor search algorithms with different molecular representations (SMILES and Morgan Fingerprints70) to sample Enamine neighboring BBs from the predicted BBs, as illustrated in colored inks in Figure 1(d). Since in SynLlama’s output, the RXN sequences are predicted concurrently with the BBs, our effective search space is constrained to Enamine BBs that can react through the specific RXN template. This smaller Enamine search space not only allows us to ensure the success rate of such forward syntheses but also allows us to effectively explore segments of the input molecule. Further details of the nearest neighbor search algorithms are provided in the Supplementary Information.
@@ 3.1 Synthesis Planning for Unseen Molecules
Having demonstrated that SynLlama models can reliably predict reaction sequences and building blocks in Table 1, we now use SynLlama to plan the synthesis of two groups of 1000 previously unseen molecules from the Enamine Diversity Set9 and the publicly available ChEMBL database8. These datasets are specific to drug-like molecules, unlike the training data used for SynLlama; the drug-related property distribution of both sets of molecules against the training data is shown in Figure S3. In this validation, we test whether known synthesizable molecules can be reconstructed accurately from the baseline and SynLlama models as summarized in Table 2.
We first consider the standard reconstruction approach used by algorithms such as SynNet54, ChemProjector55, and Synformer56 to create target molecules or analogs using BBs exclusively from the Enamine library. In this comparison, SynNet54 and ChemProjector55 serve as a baseline comparison for synthesizable chemical space coverage for RXN 1, whereas Synformer56 is the baseline comparison on the expanded RXN 2 templates. As seen in the first column of Table 2 and Supplementary Table S3, when trained with their respective reaction sets and using only Enamine BBs, SynLlama outperforms all three methods while reducing the number of training data by 40-to 60-fold.
@@
Table 2: Comparison of synthesis planning performance among different methods. Both the Enamine Diversity Set and ChEMBL Data are comprised of 1000 unseen molecules each. Details of each benchmark are described in the main text. The Morgan similarity scores include all analog molecules with successful synthesis pathways, as well as successfully reconstructed target molecules. The number of total training data and reaction set each method used for is: SynNet54 (200K, RXN 1) ChemProjector55 (128M, RXN 1), Synformer56 (85M, RXN 2), and SynLlama (2M, RXN 1 & 2). Further details are provided in Supplementary Tables S4 and S5, and identification of purchasable New $B$s with Molport69 are described in the Supplementary Information.
However, unlike baseline methods that only generate molecules using Enamine BBs, SynLlama has the extra capacity of reconstruct target molecules with commercially available BBs beyond Enamine due to its generative capabilities, even without specific training for this purpose. As seen in Table 2 the ‘New BBs’, restricted to those purchasable through Molport69, add possible synthetic pathways to reconstruct target molecules in all datasets and RXN templates. This also helps the ChEMBL data as well given that the molecules generated with the new BBs remain drug-like (Supplementary Figure S5). Since some target molecules can be synthesized through multiple pathways, either using only Enamine BBs or with the addition of New BBs, the ‘Total’ column in Table 2 reflects the number of unique target molecules reconstructed with SynLlama. With these New BBs, SynLlama’s best reconstruction rates increase to 74.1% for the 1000 molecules in the Enamine Diversity Set, and encouragingly to 28.7% for the ChEMBL data. These results show that SynLlama learns reaction chemistry such that it can predict novel BBs to increase synthetic accessibility.
When the target molecule cannot be reconstructed, we assess the quality of the analog using a molecular similarity score between the target molecule and its most similar analog using Tanimoto similarity based on 4096-bit Morgan fingerprints70. In Table 2, the similarity metrics reported are average values of all generated molecules, including target molecules that are fully reconstructed (with a score of 1). Tables S6 and S7 also provide similarity metrics based on the 4096-bit Morgan fingerprints of Murcko scaffolds71, and Gobbi 2D pharmacophore fingerprints72, while including or excluding target molecules that are fully reconstructed. Overall, these results collectively show that SynLlama is highly capable of planning synthesis for related analog molecules with very good similarity, aided most by increased synthetic pathways using purchasable building blocks.

### v5

@@ 1 Introduction
Chemical space is enormous, built up via the exponential rise in functional group combinatorics that define an increasing diverse set of molecules. Traditional approaches that design synthetic pathways of unseen molecules under well-controlled laboratory conditions have been made possible by decades of research in synthetic chemistry, as well as mechanistic studies of key reaction steps and reaction classificationcorey_general_1967, ihlenfeldt_computerassisted_1996. Using the wealth of data accumulated in libraries of chemical reactions, expert systemscorey_computer-assisted_1985, Kowalik2012 have been developed that deploy this knowledge to construct multi-step pathways to specified end-products. Such methods have become a key tool for the bench chemist, as illustrated by the Chematica software and and its follow on commercial product now known as SynthiaKowalik2012.
@@
With recent advances in artificial intelligence and deep learning, generative models have begun to contribute to enumerating molecules at the stoichiometric scale. After training on databases containing various small molecules representationspdbbind_2004, bindingdb1, irwin_zinc_2012, gaulton_chembl_2012, enamine_bb, string-based 1D generative models and structure-aware 3D de novo methods have paved the way for quick exploration of greater swaths of unseen chemical spacechithrananda_chemberta_2020, li_molbert_2021, mol-gen-gpt0, eckmann_limo_2022, wang_cmolgpt_2023, mol-gen-bg2, mol-gen-bg3, mol-gen-bg4, denovo_Guan2023TargetDiff, iminer, denovo_li2023ls, denovo_Luo2021SBDD, denovo_pocket2mol, denovo_zhang2022novo, denovo_Qian2022AlphaDrug, denovo_schneuing2022structure, denovo_zhang2023resgen, cavanagh_smileyllama_2024. However, even with their exceptional generative capabilities, these models still face one major challenge: their proposed de novo molecules lack practical guarantees of synthesizability, which limits their utility in practicesumita_hunting_2018, zhavoronkov_deep_2019, gao_synthesizability_2020. For generative approaches in drug and materials discovery to fulfill their potential, ensuring synthetic feasibility is essential to bridge the gap between in silico molecule design and the realistic applicability of computationally generated molecules.
@@
In recognition of this dissonance, efforts have been made to address the problem of poor synthesizability of de-novo-generated molecules. One line of research focuses on integrating empirical or deep-learning scoring functionsertl_estimation_sa_score_2009, coley_scscore_2018, vorsilak_syba_2020, thakkar_retrosynthetic_2021, wang_deepsa_2023, kim_dfrscore_2024, neeser_fsscore_2024, such as the synthetic accessibility (SA) score ertl_estimation_sa_score_2009 and DeepSA scorewang_deepsa_2023, into the objective functions of learning algorithms. However, optimizing only the synthesizability score can still lead to the generation of unsynthesizable molecules because the scoring functions rely on identifying common fragments or reactive centers in moleculesbilodeau_generative_2022. In addition, they often assign bad scores to complex yet synthesizable molecules that require multi-step synthetic pathways, causing generative models to miss viable candidates skoraczynski_critical_2023. Others have proposed improving synthesizability by building molecules from common molecular fragmentspodda_deep_2020, yang2021hit, chen_deep_2021, seo_molecular_2023, but they still don’t guarantee synthesis as these methods do not explicitly consider the reaction pathways to build the molecular candidates. Another line of research integrates the explicit use of computer-assisted synthetic planning (CASP) softwaregenheden_aizynthfinder_2020, tu_askcos_2025 into the optimizationguo_directly_2025. However, the computational overhead is significant and the quality of the optimized molecules can be variable.Shen2021.
@@
Alternatively, proposing synthesizable molecular candidates using commercially available building blocks and commonly known organic reaction templatescoley_robotic_2019, genheden_aizynthfinder_2020 offers better synthetic tractability over simple molecule scoring. Importantly, this strategy is appealing to bench and medicinal chemists, since it offers actionable synthesis pathways for them to examine, refine, and execute. Some recent models in this direction apply rule-based synthesis and optimization on building blocks or entire synthetic pathways to generate novel molecules with desired chemical propertiesKowalik2012, li_deep_2024, swanson_generative_2024, cretu2025synflownet, seo_generative_2024, wang_clickgen_2024. Other models condition on input molecules to propose synthetic pathways using commercially available building blocks and well-validated reaction templates for either full construction of the target molecule or the generation of structurally similar analogs in a forward synthesis manner within the predefined chemical search space.gao_amortized_2022, luo_projecting_2024, gao_generative_2024 For example, SynNetgao_amortized_2022 constructs synthetic trees via Markov Decision Processes (MDPs) and uses multilayer perceptrons to choose the next action space. More recent models such as ChemProjectorluo_projecting_2024 and Synformergao_generative_2024 use transformers to decode for the next action space and have achieved good empirical performances for target and analog molecule reconstruction.
@@
A compelling alternative is the use of Large Language Models (LLMs) due to their foundational nature and adaptability to downstream tasks. bommasani2022opportunitiesrisksfoundationmodels LLMs inherently possess extensive chemical knowledge, and recent advancements have focused on extracting and applying this knowledge for predictive and optimization tasks using natural language guidanceboiko_autonomous_2023, m_bran_augmenting_2024, yu2024llasmol, ramos_review_2024. Furthermore, after fine-tuning, LLM models can perform as good or better than chemical language models trained solely on chemical representations, all while requiring less data.cavanagh_smileyllama_2024 The efficiency and unexpected performance gains from fine-tuning LLMs thus motivates us to explore their potential in more complex tasks, such as synthesis planning, which could pave the way for new chemical discoveries.
@@
Herein, we present SynLlama, an LLM-based tool built on the open-source Llama-3.1-8B and Llama-3.2-1B foundation models  dubey_llama_2024 to deduce synthetic routes for target molecules or structurally related analogs. Specifically, the LLM component of SynLlama operates as a constrained retrosynthesis module that breaks input molecules into building blocks (BBs) via well-validated (RXN) sequences, and the reconstruction module searches commercially available BBs based on LLM predictions and builds up molecules within a diverse yet synthesizable chemical space. As an illustration of utility, SynLlama demonstrates competitive performance in key tasks for drug discovery, including synthesis planning for target and analog molecules of pharmaceutical interest and expansion around existing molecular drug hits and leads. Moreover, because of its generative nature, the LLM component of SynLlama has the added ability to explore commercially available building blocks beyond the predefined synthetic space introduced during training - an ability that previous models lack. By integrating molecular design with synthetic feasibility, SynLlama represents a step forward in bridging computational chemistry with synthetic chemistry, providing chemists with actionable and experimentally accessible molecular candidates.
@@ 2 Methods
The SynLlama workflow, illustrated in Figure 1, is designed to generate synthesizable compounds within an expanded chemical space. When an input molecule passes through this workflow, it can either be fully reconstructed through valid synthetic pathways, or the workflow will produce a structurally similar yet synthesizable analog along with its synthesis route. To transform general-purpose LLMs, like the Llama 3 models dubey_llama_2024, into expert models for synthetic pathways, we use three key components: 1) a reliable and diverse set of reaction data that covers a large synthesizable chemical space, 2) an efficient supervised fine-tuning (SFT) strategy to train a general-purpose LLM on these reaction data, and 3) a reconstruction algorithm that can convert the output of the fine-tuned LLM into valid synthesis routes, ensuring the proposed molecules lie within a commercially available chemical search space. These components are crucial for leveraging LLMs, which are known to perform well in diverse chemistry tasks hendrycks2020mmlu, wang2024mmlupro, to specialize in synthetic modeling.
@@ 2.1 Reaction Data for Training and Testing Sets
As illustrated in Figure 1(a), our defined chemical space for training consists of molecules that can be synthesized in at most five steps with Enamine building blocksenamine_bb (BBs) and 2 sets of well-validated common organic reactions (RXNs). To define the training and testing BB data, we apply a time split whereby all Enamine BBs from the August 2024 release serve as the training BBs, and all new BBs from their February 2025 release that were not in the training set comprise the testing BBs. This procedure results in $\sim$230,000 BBs for training and $\sim$13,000 BBs for testing. Later in Section 3.1 we consider the target reconstruction for unseen molecules which are taken from the Enamine Diversity Setenamine_bb and ChEMBL datasetgaulton_chembl_2012.
We define two sets of reaction templates (RXN) which operate on the Enamine BBs. RXN 1 is formulated as a set of 91 reaction templates selected by Gao et al.gao_amortized_2022 from the works of Hartenfeller et al.hartenfeller_dogs_2012 and Button et al.button_automated_2019. RXN 2 is comprised of 115 reactions selected by Gao et al.gao_generative_2024 that contains reactions used to create the Enamine REAL spaceenamine_bb plus some reactions from RXN 1. All reaction templates are accessible to both sets of BBs, thus defining the training and testing chemical spaces. As a result, there are $\sim$$10^{30}$ molecules within this space that can be represented by a synthesis path that comprises a sequence of BBs and RXNs.
@@ 2.2 Supervised Fine-tuning and Inference from SynLlama
To create the SynLlama model, we need to establish data generation protocols for supervised fine-tuning (SFT) of the Llama 3 models as schematically shown in Figure 1(c). When generating reaction data in text format, we choose to represent the BBs and intermediates along the synthetic pathway using SMILESSMILES strings, while RXNs are explicitly defined in the SMARTSSMARTS format. These structured chemical notations are designed to enhance SynLlama’s ability to systematically identify and deconstruct bonds according to RXN templates, effectively dismantling input molecules into building-block-sized fragments.
@@ 2.3 SynLlama Model Benchmarks
In Table 1, all four trained SynLlama models are evaluated on both in-distribution training data and out-of-distribution testing and ChEMBL gaulton_chembl_2012 data to assess the benchmarks outlined above. In the instruction-following benchmarks, most models exhibit strong adherence (over 90 percent) to the fine-tuned response structure across all datasets. This impressive performance indicates that fine-tuning effectively retains the specified output structure when trained with over 100,000 samples. Furthermore, all four models successfully memorized the provided RXN templates and selected the building blocks (BBs) from all predicted reactants over 99 percent of the time. This capability further enhances the coupling effectiveness of the downstream reconstruction algorithm with the SynLlama raw output, as it only requires information about reaction sequences and predicted building blocks.
In the reaction chemistry benchmarking results, a clearer trend emerges: models, regardless of their parameter size, show improved comprehension of reaction chemistry in all three datasets as the amount of training data increases. Notably, most models maintain their performance from training to testing data, but exhibit a greater decline in “Matched Reactants” and “Good Products” performance when generalizing to the ChEMBL data. The reason behind this is that the testing data are generated in the same manner as the training data but with a different set of building blocks, while the ChEMBL data occupies a different chemical space, as previously noted by Gao et al. gao_amortized_2022. Despite the reductions in their performance for ChEMBL molecules, as shown in Supplementary Figure S2, SynLlama-8B-500k and SynLlama-1B-2M can still generate complete and valid syntheses over 50% of the time without any downstream processing. These results indicates that SynLlama’s raw results alone have potential utility for synthesis planning for unseen drug-like molecules.
@@ 2.4 Reconstruction from Predicted Retrosynthesis
Using the predicted sequence of RXNs and BBs from SynLlama responses, we can synthesize the proposed target molecule or close analogs by applying the predicted reaction templates to the BBs in the inferred order, as shown in black ink in Figure 1(d). In some cases, the predicted BBs match known Enamine BBs, ensuring that the resulting molecules remain within an established chemical space for synthesis. However, due to SynLlama’s generative nature, some predicted BBs are novel while still providing valid synthesis pathways, and we only report new BBs that can be purchased from other suppliers identified by Molport molport. Therefore, while SynLlama primarily produces molecules within the predefined chemical space using Enamine BBs, its output also offers an alternative strategy for molecule construction. We will revisit this point in the Results section.
When the input molecule cannot be fully reconstructed, we generate analogs by mapping the predicted BBs from SynLlama to known Enamine BBs, thereby sampling molecules from the well-defined Enamine chemical space. Under this scenario, we use nearest neighbor search algorithms with different molecular representations (SMILES and Morgan Fingerprintsmorgan1965unique) to sample Enamine neighboring BBs from the predicted BBs, as illustrated in colored inks in Figure 1(d). Since in SynLlama’s output, the RXN sequences are predicted concurrently with the BBs, our effective search space is constrained to Enamine BBs that can react through the specific RXN template. This smaller Enamine search space not only allows us to ensure the success rate of such forward syntheses but also allows us to effectively explore segments of the input molecule. Further details of the nearest neighbor search algorithms are provided in the Supplementary Information.
@@ 3.1 Synthesis Planning for Unseen Molecules
Having demonstrated that SynLlama models can reliably predict reaction sequences and building blocks in Table 1, we now use SynLlama to plan the synthesis of two groups of 1000 previously unseen molecules from the Enamine Diversity Setenamine_bb and the publicly available ChEMBL databasegaulton_chembl_2012. These datasets are specific to drug-like molecules, unlike the training data used for SynLlama; the drug-related property distribution of both sets of molecules against the training data is shown in Figure S3. In this validation, we test whether known synthesizable molecules can be reconstructed accurately from the baseline and SynLlama models as summarized in Table 2.
We first consider the standard reconstruction approach used by algorithms such as SynNetgao_amortized_2022, ChemProjectorluo_projecting_2024, and Synformergao_generative_2024 to create target molecules or analogs using BBs exclusively from the Enamine library. In this comparison, SynNetgao_amortized_2022 and ChemProjectorluo_projecting_2024 serve as a baseline comparison for synthesizable chemical space coverage for RXN 1, whereas Synformergao_generative_2024 is the baseline comparison on the expanded RXN 2 templates. As seen in the first column of Table 2 and Supplementary Table S3, when trained with their respective reaction sets and using only Enamine BBs, SynLlama outperforms all three methods while reducing the number of training data by 40-to 60-fold.
@@
Table 2: Comparison of synthesis planning performance among different methods. Both the Enamine Diversity Set and ChEMBL Data are comprised of 1000 unseen molecules each. Details of each benchmark are described in the main text. The Morgan similarity scores include all analog molecules with successful synthesis pathways, as well as successfully reconstructed target molecules. The number of total training data and reaction set each method used for is: SynNetgao_amortized_2022 (200K, RXN 1) ChemProjectorluo_projecting_2024 (128M, RXN 1), Synformergao_generative_2024 (85M, RXN 2), and SynLlama (2M, RXN 1 & 2). Further details are provided in Supplementary Tables S4 and S5, and identification of purchasable New $B$s with Molportmolport are described in the Supplementary Information.
However, unlike baseline methods that only generate molecules using Enamine BBs, SynLlama has the extra capacity of reconstruct target molecules with commercially available BBs beyond Enamine due to its generative capabilities, even without specific training for this purpose. As seen in Table 2 the ‘New BBs’, restricted to those purchasable through Molportmolport, add possible synthetic pathways to reconstruct target molecules in all datasets and RXN templates. This also helps the ChEMBL data as well given that the molecules generated with the new BBs remain drug-like (Supplementary Figure S5). Since some target molecules can be synthesized through multiple pathways, either using only Enamine BBs or with the addition of New BBs, the ‘Total’ column in Table 2 reflects the number of unique target molecules reconstructed with SynLlama. With these New BBs, SynLlama’s best reconstruction rates increase to 74.1% for the 1000 molecules in the Enamine Diversity Set, and encouragingly to 28.7% for the ChEMBL data. These results show that SynLlama learns reaction chemistry such that it can predict novel BBs to increase synthetic accessibility.
When the target molecule cannot be reconstructed, we assess the quality of the analog using a molecular similarity score between the target molecule and its most similar analog using Tanimoto similarity based on 4096-bit Morgan fingerprintsmorgan1965unique. In Table 2, the similarity metrics reported are average values of all generated molecules, including target molecules that are fully reconstructed (with a score of 1). Tables S6 and S7 also provide similarity metrics based on the 4096-bit Morgan fingerprints of Murcko scaffoldsmurcko_scaffold, and Gobbi 2D pharmacophore fingerprintsgobbi_1998, while including or excluding target molecules that are fully reconstructed. Overall, these results collectively show that SynLlama is highly capable of planning synthesis for related analog molecules with very good similarity, aided most by increased synthetic pathways using purchasable building blocks.